OpenStudy (vulpixshay):

Mike started a savings account by depositing $9. Each month, he deposits more money than the month before. At the end of 41 months, he has saved $9,389.00. How much more does he deposit each month? $11.00 $11.50 $12.00 $12.50

OpenStudy (prathamesh_m):

Let the extra amount he deposits be x. You can consider this as an arithmetic progression where 9 is the first term and x is the common difference between successive terms. So the total amount($9389) would be equal to : 9 + (9+x) + (9+x+x) ...... = 9 + (9+x) + (9+2x) .... He deposits these amounts for 41 months so there are 41 terms in that sequence. So now you can form an equation: \[\frac{ 41 }{ 2 }([2 \times 9] + (41-1)x) = 9389\]

OpenStudy (prathamesh_m):

You can solve that equation to find x. The left hand side of that equation comes from the formula for the sum of n terms of an arithmetic progression: \[S _{n} = \frac{ n }{ 2 }(2a + (n-1)d)\]

OpenStudy (vulpixshay):

@Prathamesh_M sorry it took me so long to reply, thank you!

OpenStudy (prathamesh_m):

Welcome!